OpenStudy (anonymous):

Suppose G is a vector field with the property that divG = 5 for 2 ≤ || r || ≤ 8 and that the flux of G through the sphere of radius 4 centered at the origin is 19π. Find the flux of G through the sphere of radius 6 centered at the origin.

OpenStudy (unklerhaukus):

Flux of G \[\Phi_G=\oint_S\mathbf{\vec G}\cdot\mathrm d\mathbf{\vec S}\]

OpenStudy (unklerhaukus):

Gauss's Divergence Theorem \[\oint_S\mathbf{\vec F}\cdot\mathrm d\mathbf{\vec S}=\iiint_V(\nabla\cdot\mathbf{\vec F})\mathrm d V\]

OpenStudy (unklerhaukus):

Surface area of a sphere\[S=4\pi r^2\] Volume of a sphere\[V = \tfrac43\pi r^3\]

OpenStudy (anonymous):

so its going to be 5 * 4/3pi (6)^3?

OpenStudy (anonymous):

They're going to be proportional.

OpenStudy (unklerhaukus):

@Ldaniel come back to this question

ganeshie8 (ganeshie8):

careful you need to subtract the volumes @Ldaniel

OpenStudy (anonymous):

@UnkleRhaukus I felt asleep

OpenStudy (anonymous):

@ganeshie8 can you help me finish up

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):



OpenStudy (anonymous):

so i need to find the flux for radius 6 and radius 4

OpenStudy (anonymous):

and get the difference?

ganeshie8 (ganeshie8):

Kindof, use divergence theorem for the region between inner sphere and outer sphere

ganeshie8 (ganeshie8):

just multiply the divergence by the volume

ganeshie8 (ganeshie8):

flux through outer sphere : \[\large 19\pi + 5 (V_6 - V_4)\]

OpenStudy (anonymous):

1440pi - 425/3pi?

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

so i got 3952pi/3

OpenStudy (anonymous):

it was mark wrong

OpenStudy (anonymous):

19pi+5(288pi-(85/3)pi)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i guess is better to type everything in wolframapha to make no mistake haha

ganeshie8 (ganeshie8):

lol agree, no point in wasting time proving ur arithmetic efficiency to yourself... let the wolfram do the donkey work where possible ;)

OpenStudy (anonymous):

cam you help me with one more thing

ganeshie8 (ganeshie8):

il try, post it